There are six items total. You are to select three of them at random, putting them in a specific order. In other words, order matters. It is a permutation question. How many different arrangements of 3 can be made this way?
Found 2 solutions by stanbon, jorel1380:
Answer by stanbon(75887)

There are six items total. You are to select three of them at random, putting them in a specific order. In other words, order matters. It is a permutation question. How many different arrangements of 3 can be made this way?
---------------
# of groups of 3 = 6C3 = (6*5*4)/(1*2*3) = 20
# of arrangements of the set you pick 3! = 6
-----------
Total # of arrangements possible:: 20*6 = 120
-------------
Cheers,
Stan H.
----------

Answer by jorel1380(3719)

If the order matters, then there are 6x5x4 different ways of selecting from 6 different objects, or 120 different ways, total. ☺☺☺☺